# GS4: Generalizable Sparse Splatting Semantic SLAM

### v2

@@ Abstract
Traditional SLAM algorithms excel at camera tracking, but typically produce incomplete and low-resolution maps that are not tightly integrated with semantics prediction. Recent work integrates Gaussian Splatting (GS) into SLAM to enable dense, photorealistic 3D mapping, yet existing GS-based SLAM methods require per-scene optimization that is slow and consumes an excessive number of Gaussians. We present GS4, the first generalizable GS-based semantic SLAM system. Compared with prior approaches, GS4 runs 10× faster, uses 10× fewer Gaussians, and achieves state-of-the-art performance across color, depth, semantic mapping and camera tracking. From an RGB-D video stream, GS4 incrementally builds and updates a set of 3D Gaussians using a feed-forward network. First, the Gaussian Prediction Model estimates a sparse set of Gaussian parameters from input frame, which integrates both color and semantic prediction with the same backbone. Then, the Gaussian Refinement Network merges new Gaussians with the existing set while avoiding redundancy. Finally, we propose to optimize GS for only 1-5 iterations that corrects drift and floaters when significant pose changes are detected. Experiments on the real-world ScanNet and ScanNet++ benchmarks demonstrate state-of-the-art semantic SLAM performance, with strong generalization capability shown through zero-shot transfer to the NYUv2 and TUM RGB-D datasets.
@@ 1 Introduction
Traditional visual SLAM systems consist of several independent components, including keypoint detection, feature matching, and bundle adjustment (Mur-Artal et al., 2015; Mur-Artal & Tardós, 2017; Campos et al., 2021). Their scene representations are typically low-resolution voxels, which limit geometric detail. Thus, although these systems generally provide accurate camera localization, they struggle to generate dense, high-quality 3D maps, which are required for robotics applications such as mobile manipulation. Recent advances in differentiable rendering (Mildenhall et al., 2020; Kerbl et al., 2023a) introduces new options for scene representation in visual SLAM. For example, neural scene representations such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) have been successfully adopted in SLAM frameworks (Sucar et al., 2021; Zhu et al., 2022b; Johari et al., 2023); however, NeRF requires hours of per-scene optimization, making it computationally expensive and forcing a trade-off between reconstruction quality and training cost.
Recently, 3D Gaussians have emerged as a powerful 3D scene representation, offering fast, differentiable, and high-quality rendering capabilities (Kerbl et al., 2023b). Leveraging these advantages, Gaussian-based representations have proven highly effective for SLAM systems (Keetha et al., 2024; Matsuki et al., 2024). However, existing approaches still rely on test-time, gradient-based optimization to estimate 3D Gaussians for each scene independently, which is computationally expensive and unsuitable for real-time applications. In addition, these methods depend on heuristic Gaussian densification and pruning strategies (Kerbl et al., 2023b), often producing overly dense representations that fail to scale to large, real-world environments.
@@
We propose GS4, the first generalizable Gaussian splatting semantic SLAM approach on monocular RGB-D sequences. Results showed that GS4 obtains state-of-the-art on real ScanNet and ScanNet++ scenes, and also zero-shot generalizes to the real NYUv2 and TUM RGB-D datasets without any fine-tuning.
@@
Our proposed few-iteration Gaussian optimization significantly improves reconstruction quality with a small additional computational cost.
@@ 2 Related Work
Traditional SLAM: Early visual SLAM methods (Mur-Artal et al., 2015) demonstrated robust localization through effective keypoint detection and matching, which resulted in sparse 3D reconstructions. While these approaches provided reliable localization, the sparse nature of the reconstructed maps limited their utility in applications requiring detailed 3D maps. To address this issue, dense visual SLAM (Kerl et al., 2013; Czarnowski et al., 2020) focused on constructing detailed maps to support applications like augmented reality (AR) and robotics. Prior methods (Canelhas et al., 2013; Dai et al., 2017b; Newcombe et al., 2011; Bylow et al., 2013; Whelan et al., 2013; Prisacariu et al., 2017) employ representations based on Signed Distance Fields (SDF), rather than relying on sparse representations such as point clouds or grids. However, these approaches often suffer from over-smoothed reconstruction, failing to capture fine details crucial for certain tasks.
@@
NeRF-based SLAM: Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) gained popularity as a 3D scene representation due to its ability to generate accurate and dense reconstructions. NeRF employs Multi-Layer Perceptron (MLP) to encode scene information and performs volume rendering by querying opacity and color along pixel rays. Methods such as iMAP (Sucar et al., 2021), NICE-SLAM (Zhu et al., 2022b), and ESLAM (Johari et al., 2023) incorporate this implicit scene representation into SLAM, leveraging NeRF’s high-fidelity reconstructions to improve both localization and mapping. DNS-SLAM (Li et al., 2023) further incorporates semantic information into the framework. However, the volumentric rendering process in NeRF is costly, often requiring trade-offs such as limiting the number of pixels during rendering, These trade-offs, while improving efficiency, may compromise the system’s accuracy in both localization and mapping.
@@
GS-based SLAM: 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023b) employs splatting rasterization instead of ray marching. This approach iterates over 3D Gaussian primitives rather than marching along rays, resulting in a more expressive and efficient representation capable of capturing high-fidelity 3D scenes with significantly faster rendering speed. Hence, GS-based SLAM systems achieve improved accuracy and speed in dense scene reconstruction. SplaTAM (Keetha et al., 2023) introduces silhouette-guided rendering to support structured map expansion, enabling efficient dense visual SLAM. Gaussian Splatting SLAM (Matsuki et al., 2024) integrates novel Gaussian insertion and pruning strategies, while GS-ICP SLAM (Ha et al., 2024) and RTG-SLAM (Peng et al., 2024) combine ICP with 3DGS to achieve both higher speed and superior map quality. Expanding upon these advancements, SGS-SLAM (Li et al., 2024b), OVO-SLAM (Martins et al., 2024), SemGauss-SLAM (Zhu et al., 2025) and GS3LAM (Li et al., 2024a) extend 3D Gaussian representations to include semantic scene understanding. However, existing GS-based SLAM methods employ per-scene optimization, requiring iterative refinement of Gaussians initialized from keyframes through rendering supervision. As a result, they all rely on additional segmentation models to predict semantic labels for each image, creating computational overhead.
Feed-forward Models for GS: Recent research has introduced feed-forward approaches for scene-level 3DGS reconstruction using generalizable models (Charatan et al., 2024; Chen et al., 2024; Liu et al., 2025). Unlike previous methods, GS-LRM (Zhang et al., 2024) avoids specialized 3D structural designs, instead using a transformer to achieve state-of-the-art results. However, to the best of our knowledge, feed-forward models have been applied on a small number of images and have not yet been introduced in GS-based semantic SLAM approaches with thousands of frames.
@@ 3.1 Gaussian Splatting
where is a quantity of a pixel to be rendered, which can be color, depth or semantic label, and is that quantity of the -th 3D Gaussian, while is its visibility, computed from opacity and covariance parameters (determined by rotation and scale). Following Keetha et al. (2024), We also render a silhouette image to determine visibility .
@@ 3.2.1 Backbone for Gaussian Prediction
We train a transformer model to regress 3D GS parameters from an image with a known camera pose (from tracking, described in Sec. 3.3.1), while simultaneously assigning semantic labels to these 3D Gaussians. We start with a pre-trained 2D image segmentation model such as Mask2Former (Cheng et al., 2022) or AutoFocusFormer (Ziwen et al., 2023), which encodes an image into encoder tokens and decoder tokens (from their image decoder) at several progressively downsampled levels , with usually. We concatenate an RGB image and a depth image resulting in a 4-channel feature map that is fed into the model:
@@
To supervise the prediction of semantic segmentation, we follow the setup in Mask2Former (Cheng et al., 2022). We denote the corresponding segmentation loss by . In addition, we render images at supervision views—comprising the current input view and randomly selected novel views that overlap with the current input—using the predicted Gaussians from the current input, and minimize RGB-D and semantic rendering loss. For novel view supervision, we focus solely on areas visible in the input view, ensuring that the optimization process focuses on regions consistently observed across both input and novel views. We explain the loss functions used during training below.
RGB Rendering Loss. Following previous work Zhang et al. (2024); Ziwen et al. (2024), we use a combination of the Mean Squared Error (MSE) loss and Perceptual loss: where is the weight for the perceptual loss.
@@ 3.2.2 Gaussian Refinement Network
The previous subsection predicts Gaussian parameters from a single frame. In our SLAM system, as new frames arrive, we insert Gaussians from the frame into unmapped regions of the current 3D reconstruction. We perform co-visibility, which involves rendering a silhouette image for the new frame to identify the regions where new Gaussians should be inserted. To ensure that the combined set of Gaussians accurately represents the scene, we propose a novel Gaussian Refinement Network to refine both the existing Gaussians in the 3D map and the newly added ones, enabling their effective merging. The input to the network includes the features and 3D positions of all Gaussians from the 3D map that are visible in the new frame, as well as Gaussians from the new frame. We process these using several local-attention transformer layers with 3D neighborhoods in the world coordinate system to fuse and update the features for each Gaussian. Subsequently, MLP layers predict updates , , and for each Gaussian. These updates refine the Gaussians to accurately render both current and previous views. To supervise the network, we render the current view along with previous overlapping views. The total training loss is:
@@ 3.3 The SLAM System
An overview of the system is summarized in Fig. 3. The system always maintains a set of 3D Gaussians representing the entire scene. For each new RGB-D image, the Gaussian prediction network predicts 3D Gaussian parameters, which can be rendered into high-fidelity color, depth, and semantic images. The Gaussian refinement network refines both the existing Gaussians in the 3D map and the newly added ones to accurately render both current and previous views. During testing, we occasionally run few-iteration test-time optimization and refine 3D Gaussians in the map to reflect camera pose updates from loop closure and bundle adjustment in the tracking module.
@@ 3.3.1 Tracking and Global Bundle Adjustment
For camera tacking in our SLAM system, we adopt a tracker used in GO-SLAM (Zhang et al., 2023) which is an enhanced version of DROID-SLAM’s tracking module (Teed & Deng, 2021). It first predicts motion in every frame. In local camera tracking, a keyframe is initialized when sufficient motion is detected, and loop closure (LC) is performed. Meanwhile, global localization performs full bundle adjustment (BA) for real-time global refinement once the system contains more than 25 keyframes. Both LC and BA help address the problem of accumulated errors and drift that can occur during the localization process.
3.3.2 Few-Iteration Gaussian Optimization
Loop closure and bundle adjustment are essential components in SLAM systems, employed to correct accumulated drift and adjust the camera poses of previous frames. However, these adjustments can cause Gaussians inserted based on earlier, uncorrected poses to misalign with the scene, leading to inaccurate rendering and mapping. It is crucial to implement a mechanism that updates the Gaussians in the 3D map following pose corrections. To address this issue, we propose using rendering-based optimization to update the Gaussian parameters , , and with only a few iterations. We render RGB-D images for the top-k frames, selected based on significant pose changes. This approach maintains the consistency of the 3D map with updated camera poses. To enhance the efficiency of this optimization, we employ the batch rendering technique from Ye et al. (2024). We omit semantic image rendering to improve system efficiency. For few-iteration optimization, we add a SSIM term to the RGB loss, following Kerbl et al. (2023b):
where is set to for all experiments.
@@ 4.1 Experimental Setup
Training Settings. We train our Gaussian prediction and refinement networks entirely on RGB-D videos from the real ScanNet datasets We exclude the six standard SLAM test scenarios and use all remaining training and validation scenes, supervising with 20 common semantic classes. For ScanNet++v2, we exclude the two SLAM test scenarios and use all other available training scenes, supervising with the 100 most common semantic classes. We adopt AutoFocusFormer (Ziwen et al., 2023) as the backbone for both Mask2Former and Gaussian prediction, using the second stage of the backbone as the prediction stage. Additionally, we experiment with Swin Transformer (Liu et al., 2021) as an alternative backbone. The detailed results are provided in the appendix. Following the low-to-high resolution curriculum of (Ziwen et al., 2024), we train the Gaussian prediction network in three stages with input resolutions of 256×256, 480×480, and 640×480. In the first two stages, images are resized such that the shorter side is 256 or 480 pixels and then center-cropped to a square. On ScanNet++v2, we use resolutions of 256x384. For the refinement network, which processes multiple consecutive frames, we adopt a progressive training schedule: beginning with two frames, then four, and finally eight.
@@
Datasets. During testing, we evaluate our method on six real-world scenes on ScanNet (Dai et al., 2017a) , and six real-world scenes on ScanNet++ (Yeshwanth et al., 2023). ScanNet++ has 11 scenes with continuous camera trajectory, but SplaTAM and SGS-SLAM completely fail in 5 scenes, respectively. Hence we only use the 6 scenes they can finish running, so that their numbers are not infinitely bad. Additionally, we perform zero-shot experiments on real scenes from NYUv2 (Nathan Silberman & Fergus, 2012) and TUM RGB-D (Sturm et al., 2012a).
@@
Metrics. We use PSNR, Depth-L1 (Zhu et al., 2022a), SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018) to evaluate the reconstruction and rendering quality. We additionally report reconstruction metrics such as Accuracy, Completion, Completion Ratio (7cm) and F1 (7cm) in the appendix. For GS-based SLAM methods, we also report the number of Gaussians. For semantic segmentation, we report the mean Intersection over Union (mIoU). To evaluate the accuracy of the camera pose, we adopt the average absolute trajectory error (ATE RMSE) (Sturm et al., 2012b).
@@
Baselines. We compare our method against several state-of-the-art approaches: NeRF-based SLAM methods, including NICE-SLAM (Zhu et al., 2022a), GO-SLAM (Zhang et al., 2023), and Point-SLAM (Sandström et al., 2023); 3D Gaussian-based SLAM methods such as SplaTAM (Keetha et al., 2024), RTG-SLAM (Peng et al., 2024), and GS-ICP SLAM (Ha et al., 2024); and semantic 3D Gaussian-based SLAM methods, including SGS-SLAM (Li et al., 2024b), GS3LAM (Li et al., 2024a) and OVO-Gaussian-SLAM (Martins et al., 2024). SGS-SLAM, GS3LAM and OVO-Gaussian-SLAM are the only semantic SLAM methods available for comparison since the code is not available for other semantic SLAM approaches. Note that SGS-SLAM (Li et al., 2024b) and GS3LAM (Li et al., 2024a) employ test-time optimization using ground truth semantic labels on the test set. SGS-SLAM and GS3LAM have been shown to outperform all other existing semantic SLAM methods (Zhu et al., 2024; Li et al., 2023). To ensure a fair comparison and simulating SLAM applications in real-world scenarios where ground truth semantic labels are unavailable, we trained a 2D segmentation model using a Swin backbone (Liu et al., 2021) with Mask2Former (Cheng et al., 2022) on ScanNet and ScanNet++, following the same training strategy as our model, and used predicted semantic labels to supervise SGS-SLAM nad GS3LAM.

### v3

@@ Abstract
Traditional SLAM algorithms excel at camera tracking, but typically produce incomplete and low-resolution maps that are not tightly integrated with semantics prediction. Recent work integrates Gaussian Splatting (GS) into SLAM to enable dense, photorealistic 3D mapping, yet existing GS-based SLAM methods require per-scene optimization that is slow and consumes an excessive number of Gaussians. We present GS4, the first generalizable GS-based semantic SLAM system. Compared with prior approaches, GS4 runs 10× faster, uses 10× fewer Gaussians, and achieves state-of-the-art performance across color, depth, semantic mapping and camera tracking. From an RGB-D video stream, GS4 incrementally builds and updates a set of 3D Gaussians using a feed-forward network. First, the Gaussian Prediction Model estimates a sparse set of Gaussian parameters from input frame, which integrates both color and semantic prediction with the same backbone. Then, the Gaussian Refinement Network merges new Gaussians with the existing set while avoiding redundancy. Finally, when significant pose changes are detected, we perform only 1–5 iterations of joint Gaussian–pose optimization to correct drift, remove floaters, and further improve tracking accuracy. Experiments on the real-world ScanNet and ScanNet++ benchmarks demonstrate state-of-the-art semantic SLAM performance, with strong generalization capability shown through zero-shot transfer to the NYUv2 and TUM RGB-D datasets.
@@ 1 Introduction
Traditional visual SLAM systems consist of several independent components, including keypoint detection, feature matching, and bundle adjustment [orb-slam, orbslam2, orb-slam3]. Their scene representations are typically low-resolution voxels, which limit geometric detail. Thus, although these systems generally provide accurate camera localization, they struggle to generate dense, high-quality 3D maps, which are required for robotics applications such as mobile manipulation. Recent advances in differentiable rendering [mildenhall2020nerf, gaussiansplatting] introduces new options for scene representation in visual SLAM. For example, neural scene representations such as Neural Radiance Fields (NeRF) [mildenhall2020nerf] have been successfully adopted in SLAM frameworks [imap, nice-slam, eslam]; however, NeRF requires hours of per-scene optimization, making it computationally expensive and forcing a trade-off between reconstruction quality and training cost.
Recently, 3D Gaussians have emerged as a powerful 3D scene representation, offering fast, differentiable, and high-quality rendering capabilities [kerbl3Dgaussians]. Leveraging these advantages, Gaussian-based representations have proven highly effective for SLAM systems [keetha2024splatam, Matsuki:Murai:etal:CVPR2024]. However, existing approaches still rely on test-time, gradient-based optimization to estimate 3D Gaussians for each scene independently, which is computationally expensive and unsuitable for real-time applications. In addition, these methods depend on heuristic Gaussian densification and pruning strategies [kerbl3Dgaussians], often producing overly dense representations that fail to scale to large, real-world environments.
@@
We propose GS4, the first generalizable Gaussian splatting semantic SLAM approach on RGB-D sequences. Results showed that GS4 obtains state-of-the-art on real ScanNet and ScanNet++ scenes, and also zero-shot generalizes to the real NYUv2 and TUM RGB-D datasets without any fine-tuning.
@@
Our proposed few-iteration joint Gaussian-pose optimization significantly improves reconstruction quality with a small additional computational cost, while also slightly improving tracking accuracy when high-quality ground-truth RGB-D data are available.
@@ 2 Related Work
Traditional SLAM: Early visual SLAM methods [orb-slam] demonstrated robust localization through effective keypoint detection and matching, which resulted in sparse 3D reconstructions. While these approaches provided reliable localization, the sparse nature of the reconstructed maps limited their utility in applications requiring detailed 3D maps. To address this issue, dense visual SLAM [kerl2013dense, czarnowski2020deepfactors] focused on constructing detailed maps to support applications like augmented reality (AR) and robotics. Prior methods [canelhas2013sdf, dai2017bundlefusion, kinectfusion, bylow2013real, whelan2013robust, prisacariu2017infinitam] employ representations based on Signed Distance Fields (SDF), rather than relying on sparse representations such as point clouds or grids. However, these approaches often suffer from over-smoothed reconstruction, failing to capture fine details crucial for certain tasks.
@@
NeRF-based SLAM: Neural Radiance Fields (NeRF) [mildenhall2020nerf] gained popularity as a 3D scene representation due to its ability to generate accurate and dense reconstructions. NeRF employs Multi-Layer Perceptron (MLP) to encode scene information and performs volume rendering by querying opacity and color along pixel rays. Methods such as iMAP [imap], NICE-SLAM [nice-slam], and ESLAM [eslam] incorporate this implicit scene representation into SLAM, leveraging NeRF’s high-fidelity reconstructions to improve both localization and mapping. DNS-SLAM [li2023dnsslamdenseneural] further incorporates semantic information into the framework. However, the volumentric rendering process in NeRF is costly, often requiring trade-offs such as limiting the number of pixels during rendering, These trade-offs, while improving efficiency, may compromise the system’s accuracy in both localization and mapping.
@@
GS-based SLAM: 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] employs splatting rasterization instead of ray marching. This approach iterates over 3D Gaussian primitives rather than marching along rays, resulting in a more expressive and efficient representation capable of capturing high-fidelity 3D scenes with significantly faster rendering speed. Hence, GS-based SLAM systems achieve improved accuracy and speed in dense scene reconstruction. SplaTAM [splatam] introduces silhouette-guided rendering to support structured map expansion, enabling efficient dense visual SLAM. Gaussian Splatting SLAM [Matsuki:Murai:etal:CVPR2024] integrates novel Gaussian insertion and pruning strategies, while GS-ICP SLAM [ha2024rgbdgsicpslam] and RTG-SLAM [peng2024rtgslam] combine ICP with 3DGS to achieve both higher speed and superior map quality. Expanding upon these advancements, SGS-SLAM [li2024sgs], OVO-SLAM [martins2024ovo], SemGauss-SLAM [zhu2025semgaussslamdensesemanticgaussian] and GS3LAM [li2024gs3lam] extend 3D Gaussian representations to include semantic scene understanding. However, existing GS-based SLAM methods employ per-scene optimization, requiring iterative refinement of Gaussians initialized from keyframes through rendering supervision. As a result, they all rely on additional segmentation models to predict semantic labels for each image, creating computational overhead.
Feed-forward Models: Recently, MV-DUSt3R+ [tang2024mv] and VGGT [wang2025vggt] have demonstrated multi-view inference by extending the DUSt3R [dust3r_cvpr24] architecture for multi-view reconstruction. Meanwhile, several works have introduced feed-forward approaches for scene-level 3DGS reconstruction using generalizable models [charatan23pixelsplat, chen2024mvsplat, liu2025mvsgaussian, gslrm2024, jiang2025anysplat, xu2024depthsplat]. In particular, pixelSplat [charatan23pixelsplat] predicts 3D Gaussians directly from image features, while DepthSplat [xu2024depthsplat] connects Gaussian splatting and depth estimation, studying their interaction to enable feed-forward 3D Gaussian reconstruction from multi-view images. AnySplat [jiang2025anysplat] further predicts both 3D Gaussian primitives and camera intrinsics/extrinsics for uncalibrated multi-view inputs. However, to the best of our knowledge, these feed-forward models have so far been applied only to a relatively small number of input images and have not been scaled to the thousands of frames typical in SLAM, nor have they been incorporated into GS-based semantic SLAM systems operating over such long sequences.
@@ 3.1 Gaussian Splatting
where is a quantity of a pixel to be rendered, which can be color, depth or semantic label, and is that quantity of the -th 3D Gaussian, while is its visibility, computed from opacity and covariance parameters (determined by rotation and scale). Following [keetha2024splatam], We also render a silhouette image to determine visibility .
@@ 3.2.1 Backbone for Gaussian Prediction
We train a transformer model to regress 3D GS parameters from an image with a known camera pose (from tracking, described in Sec. 3.3.1), while simultaneously assigning semantic labels to these 3D Gaussians. We start with a pre-trained 2D image segmentation model such as Mask2Former [cheng2021mask2former] or AutoFocusFormer [autofocusformer], which encodes an image into encoder tokens and decoder tokens (from their image decoder) at several progressively downsampled levels , with usually. We concatenate an RGB image and a depth image resulting in a 4-channel feature map that is fed into the model:
@@
To supervise the prediction of semantic segmentation, we follow the setup in Mask2Former [cheng2021mask2former]. We denote the corresponding segmentation loss by . In addition, we render images at supervision views—comprising the current input view and randomly selected novel views that overlap with the current input—using the predicted Gaussians from the current input, and minimize RGB-D and semantic rendering loss. For novel view supervision, we focus solely on areas visible in the input view, ensuring that the optimization process focuses on regions consistently observed across both input and novel views. We explain the loss functions used during training below.
RGB Rendering Loss. Following previous work [gslrm2024, ziwen2024llrm], we use a combination of the Mean Squared Error (MSE) loss and Perceptual loss: where is the weight for the perceptual loss.
@@ 3.2.2 Gaussian Refinement Network
The previous subsection predicts Gaussian parameters from a single frame. In our SLAM system, as new frames arrive, we insert Gaussians from the frame into unmapped regions of the current 3D reconstruction. We perform co-visibility, which involves rendering a silhouette image for the new frame to identify the regions where new Gaussians should be inserted [splatam]. To ensure that the combined set of Gaussians accurately represents the scene, we propose a novel Gaussian Refinement Network to refine both the existing Gaussians in the 3D map and the newly added ones, enabling their effective merging. The input to the network includes the features and 3D positions of all Gaussians from the 3D map that are visible in the new frame, as well as Gaussians from the new frame. We process these using several local-attention transformer layers with 3D neighborhoods in the world coordinate system to fuse and update the features for each Gaussian. Subsequently, MLP layers predict updates , , and for each Gaussian. These updates refine the Gaussians to accurately render both current and previous views. To supervise the network, we render the current view along with previous overlapping views. The total training loss is:
@@ 3.3 The SLAM System
An overview of the system is summarized in Fig. 3. The system always maintains a set of 3D Gaussians representing the entire scene. For each new RGB-D image, the Gaussian prediction network predicts 3D Gaussian parameters, which can be rendered into high-fidelity color, depth, and semantic images. The Gaussian refinement network refines both the existing Gaussians in the 3D map and the newly added ones to accurately render both current and previous views. During testing, we occasionally run few-iteration test-time optimization and refine 3D Gaussians in the map to reflect camera pose updates from loop closure and bundle adjustment in the tracking module, while jointly optimizing the camera poses of these frames.
@@ 3.3.1 Tracking and Global Bundle Adjustment
Our main contribution is on the mapping side, hence for camera tracking in our SLAM system, we can adopt any off-the-shelf algorithm [zhang2023goslam, murai2025mast3r]. A SLAM tracking system usually consists of two components: local camera tracking and global localization. In local camera tracking, a keyframe is initialized when sufficient motion is detected, and loop closure (LC) is performed. Meanwhile, global localization performs full bundle adjustment (BA) for real-time global refinement once the system contains more than 25 keyframes. Both LC and BA help address the problem of accumulated errors and drift that can occur during the localization process.
3.3.2 Few-Iteration Joint GaussianPose Optimization
Loop closure and bundle adjustment are essential components in SLAM systems, employed to correct accumulated drift and adjust the camera poses of previous frames. However, these adjustments can cause Gaussians inserted based on earlier, uncorrected poses to misalign with the scene, leading to inaccurate rendering and mapping. It is crucial to implement a mechanism that updates the Gaussians in the 3D map following pose corrections. To address this issue, we propose using rendering-based optimization to update the Gaussian parameters , , , and with only a few iterations. Specifically, we re-render RGB-D images for the top- frames, selected based on significant pose changes, and jointly minimize a rendering loss with respect to both the Gaussian parameters and the camera poses of these frames. The optimized camera poses are then fed back into the tracking system, and used for loop closure and bundle adjustment. This joint optimization keeps the 3D map consistent with the corrected poses while also leveraging the full 3D map to further improve pose accuracy. To enhance the efficiency of this optimization, we employ the batch rendering technique from [ye2024gsplatopensourcelibrarygaussian]. We omit semantic image rendering to improve system efficiency. For few-iteration optimization, we add a SSIM term to the RGB loss, following [kerbl3Dgaussians]:
where is set to for all experiments. We perform 1–5 iterations in our experiments.
@@ 4.1 Experimental Setup
Training Settings. We train our Gaussian prediction and refinement networks entirely on RGB-D videos from the real ScanNet datasets. We exclude the six standard SLAM test scenarios and use all remaining training and validation scenes, supervising with 20 common semantic classes. For ScanNet++, we exclude the SLAM test scenarios and use all other available training scenes, supervising with the 100 most common semantic classes. We adopt AutoFocusFormer [autofocusformer] as the backbone for both Mask2Former and Gaussian prediction, using the second stage of the backbone as the prediction stage. Following the low-to-high resolution curriculum of [ziwen2024llrm], we train the Gaussian prediction network in three stages with input resolutions of 256×256, 480×480, and 640×480. In the first two stages, images are resized such that the shorter side is 256 or 480 pixels and then center-cropped to a square. On ScanNet++, we use resolutions of 256x384. For the refinement network, which processes multiple consecutive frames, we adopt a progressive training schedule: beginning with two frames, then four, and finally eight.
@@
Evaluation Datasets and Settings. During testing, we evaluate our method on six real-world scenes from ScanNet [dai2017scannet], which are commonly used as SLAM benchmarks, and six real-world scenes from ScanNet++ [yeshwanth2023scannet++]. ScanNet++ validation set contains 11 scenes with continuous camera trajectories, but SplaTAM and SGS-SLAM completely fail on 5 of them. Hence we only use the 6 scenes they can finish running, so that their numbers are not infinitely bad. Note that our approach can successfully run on all 11 scenes. Additionally, we perform zero-shot experiments on real scenes from NYUv2 [Silberman:ECCV12] and TUM RGB-D [sturm12iros]. For ScanNet, NYUv2, and TUM RGB-D, we evaluate rendering performance on every 5th frame of each sequence. For ScanNet++, we evaluate all training views of each RGB-D video and additionally hold out the novel views provided by the dataset.
@@
Evaluation Metrics. We use PSNR, Depth-L1 [Zhu2022CVPR], SSIM [1284395], and LPIPS [zhang2018perceptual] to evaluate the reconstruction and rendering quality. We additionally report reconstruction metrics such as Accuracy, Completion, Completion Ratio (7cm) and F1 (7cm) in the appendix. For GS-based SLAM methods, we also report the number of Gaussians. For semantic segmentation, we report the mean Intersection over Union (mIoU). To evaluate the accuracy of the camera pose, we adopt the average absolute trajectory error (ATE RMSE) [6385773].
@@
Baselines. We compare our method against several state-of-the-art approaches: NeRF-based SLAM methods, including NICE-SLAM [Zhu2022CVPR], GO-SLAM [zhang2023goslam], and Point-SLAM [Sandström2023ICCV]; 3D Gaussian-based SLAM methods such as SplaTAM [keetha2024splatam], RTG-SLAM [peng2024rtgslam], and GS-ICP SLAM [ha2024rgbdgsicpslam]; and semantic 3D Gaussian-based SLAM methods, including SGS-SLAM [li2024sgs], GS3LAM [li2024gs3lam] and OVO-Gaussian-SLAM [martins2024ovo]. SGS-SLAM, GS3LAM and OVO-Gaussian-SLAM are the only semantic SLAM methods available for comparison since the code is not available for other semantic SLAM approaches. Note that SGS-SLAM [li2024sgs] and GS3LAM [li2024gs3lam] employ test-time optimization using ground truth semantic labels on the test set. SGS-SLAM and GS3LAM have been shown to outperform all other existing semantic SLAM methods [zhu2024sni, li2023dnsslamdenseneural]. To ensure a fair comparison and simulating SLAM applications in real-world scenarios where ground truth semantic labels are unavailable, we trained a 2D segmentation model using a Swin backbone [liu2021Swin] with Mask2Former [cheng2021mask2former] on ScanNet and ScanNet++, following the same training strategy as our model, and used predicted semantic labels to supervise SGS-SLAM nad GS3LAM.